The mass of a sphere, M, is directly proportional to r cubed, where r is the radius of the sphere in cm

When r = 10 cm, M = 200 grams.


Work out the mass of a sphere, made of the same material, with a radius of 15 cm


1
Expert's answer
2021-01-29T05:03:38-0500
"M=Cr^3, C=const"


Solve for "C"


"C=\dfrac{M}{r^3}"

Given "M=200\ g," when "r=10\ cm"


"C=\dfrac{200g}{(10cm)^3}=0.2\ g/cm^3"

When "r=15\ cm"


"M=0.2\ g/cm^3\cdot(15cm)^3=675g"

When r = 15 cm, M = 675 grams.